Please help me solve the following equations
32.) x+7y=5 (1)
x=9-7y (2)

Answer by mananth(16946)

x+7y=5 (1)
x=9-7y (2)

x+7y=9, y=-x/7+9/7
x+7y=5, y=-x/7 +5/7

subtract the equations
we get 0=4 (absurd)
The lines are having the same slope and different y intercepts.
Hence they are parallel lines.
NO solution